OpenStudy (anonymous):

Solve for x: 2|x − 3| + 1 = 7

OpenStudy (anonymous):

x = 0, x = 6 x = −1, x = 6 x = 0, x = 9 No solutions

jimthompson5910 (jim_thompson5910):

2|x - 3| + 1 = 7 2|x - 3| = 7-1 2|x - 3| = 6 |x - 3| = 6/2 |x - 3| = 3 x - 3 = 3 or x - 3 = -3 I'll let you finish

OpenStudy (anonymous):

x - 3 = -3 +3 +3 x=0 No Solution??

jimthompson5910 (jim_thompson5910):

0 is a solution, so it's NOT "No solution"

OpenStudy (anonymous):

ok would it be a then???

jimthompson5910 (jim_thompson5910):

correct, it is A

OpenStudy (anonymous):

ok thanks

jimthompson5910 (jim_thompson5910):

np